OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)^2 + (y – 1)^2 = 64 (x – 2)^2 + (y + 1)^2 = 64 (x – 2)^2 + (y + 1)^2 = 8 (x + 2)^2 + (y – 1)^2 = 8

zepdrix (zepdrix):

Equation of a circle: \[\large(x-h)^2+(y-k)^2=r^2\] So our radius is 8. Looking at the equation, what should r^2 be?

zepdrix (zepdrix):

This equation is for a circle centered at (h,k), sorry I forgot to mention that. :)

OpenStudy (anonymous):

sorry, i'm still stuck on whether it's C or D /:

zepdrix (zepdrix):

We can immediately determine that it is neither C nor D. If our Radius = 8 Then our Radius Squared (in the equation) = 64 yes?

OpenStudy (anonymous):

ohh

zepdrix (zepdrix):

\[(x-h)^2+(y-k)^2=r^2\] Plugging in our center points and our radius, gives us this: \[(x-2)^2+(y-(-1))^2=8^2\] The negative signs inside the brackets might be a little confusing. The equation is defined with -h and -k, so we should get the opposite of our coordinates inside the brackets. Make sense, or still confused? :)

OpenStudy (anonymous):

would it be (x – 2)^2 + (y + 1)^2 = 64 then?

zepdrix (zepdrix):

Yes, looks good! :)

OpenStudy (anonymous):

thank you so much! :)